# Fair Concurrent Training of Multiple Models in Federated Learning

### v1

@@ IV-B Convergence Guarantees
Let $\overline{f}_{s}(\alpha)=\frac{f_{s}\left(\textbf{w}_{s}(t+1)\right)^{\alpha}}{\sum_{s^{\prime}}f_{s^{\prime}}\left(\textbf{w}_{s}(t+1)\right)^{\alpha}}$. When $p_{k}=\frac{1}{K}$, we can rewrite $\sum_{Sel\in\mathbb{S}(t)}\sum_{k\in Sel}B_{Sel}^{s}(\alpha)\left(\frac{p_{k}}{\sum_{k^{\prime}\in S(t)}p_{k’}}\right)^{2}$ as $\sum_{j=1}^{K}\frac{1}{j}\left(\overline{f}_{s}(\alpha)\right)^{j}\left(1-\overline{f}_{s}(\alpha)\right)^{K-j}{K\choose j}$.
Since $\sum_{j=1}^{K}\left(\overline{f}_{s}(\alpha)\right)^{j}\left(1-\overline{f}_{s}(\alpha)\right)^{K-j}{K\choose j}=1$, we can view this as a weighted average of $1,\frac{1}{2},\ldots,\frac{1}{K}$. It then suffices to show that, as $\alpha$ increases, we place more weight on smaller terms. Note that $\frac{d}{d\alpha}\left(\left(\overline{f}_{s}(\alpha)\right)^{j}\left(1-\overline{f}_{s}(\alpha)\right)^{K-j}\right)\propto\left(\frac{d\overline{f}_{s}(\alpha)}{\alpha}\right)\left(j-K\overline{f}_{s}(\alpha)\right).$ It then suffices to show that $\frac{d\overline{f}_{s}(\alpha)}{\alpha}>0$; if so, the terms with larger $j$ will increase as $\alpha$ increases. Hence, $\frac{d\overline{f}_{s}(\alpha)}{\alpha}\propto\sum_{s^{\prime}=1}^{S}f_{s^{\prime}}^{\alpha}\left(\log f_{s}-\log f_{s^{\prime}}\right)>0$ as we have assumed task $s$ has the highest loss $f_{s}$. ∎
As the fairness parameter $\alpha$ increases, our client-task allocation algorithm ensures that the worst-performing task makes more progress towards its training goal, thus helping to equalize the tasks’ performance and fulfilling the intuition of a higher $\alpha$ leading to more fair outcomes. The result further suggests that an alternative client-task allocation algorithm can explicitly enforce fairness across the upper-bounds in (4). However, this approach requires integer programming methods and estimating unknown constants like $\sigma^{2}$ for each task, making it less practical than FedFairMMFL. Finally, we provide the convergence bound for FedFairMMFL:
@@
With Assumptions 1-4, under a decaying learning rate $\eta_{t}=\frac{1}{\mu(t+\gamma)}$ and under our difficulty-aware client-task allocation algorithm, the error after $T$ iterations of federated learning for a specific task $s$ in multiple-model federated learning is
@@
Proof Sketch. We follow the outline of the proof of [22]. As our client selection scheme is different from that in [22], we use the upper bound of $\mathbb{E}\left[\|\overline{\textbf{w}}^{(t+1)}_{s}-\textbf{w}_{s}^{*}\|^{2}\right]$ from Theorem 4, which incorporates our client-task allocation scheme. Next, we upper bound $\Delta_{t+1}=\sum_{Sel\in\mathbb{S}}B_{Sel(t)}^{s}(\alpha)P^{k,2}_{Sel}$ by $\frac{1}{2}$. Firstly, the inner summation $P^{k,2}_{S(t)}$ can be re-written as $\sum_{k\in Sel}\sum_{k^{\prime}\in Sel,k^{\prime}\neq k}\frac{p_{k}}{(\sum_{k^{{}^{\prime\prime}}\in Sel}p_{k^{\prime\prime}})}\frac{p_{k^{\prime}}}{(\sum_{k^{{}^{\prime\prime}}\in Sel}p_{k^{\prime\prime}})}=\frac{(\sum_{k^{{}^{\prime\prime}}\in Sel}p_{k^{\prime\prime}})^{2}-\sum_{k\in Sel}p_{k}^{2}}{2(\sum_{k^{{}^{\prime\prime}}\in Sel}p_{k^{\prime\prime}})^{2}}\leq\frac{1}{2}.$ Secondly, the outer summation $\sum_{Sel}B_{Sel}^{s}(\alpha(t))$ is a binomial distribution with binomial parameter $f_{s}^{\beta}(t)/{\sum_{s^{\prime}\in S}f_{s^{\prime}}^{\beta}(t)}$, and this sums to 1. Now we have a bound of $\Delta_{t+1}$ in terms of $\Delta_{t}$ and the learning rates $\eta_{t}$. Next, we use induction to bound $\Delta_{t+1}$ in terms of time independent variables. Finally, we use the L-smoothness of $F_{s}$ to obtain the result of this corollary.
@@ V Client Incentive-Aware Fairness across Tasks
In this section, we account for the fact that clients may have different willingness to train each model. Prior incentive-aware FL work [31] aimed to maximize the total number of users performing either one (or many) tasks. Here, we would like to ensure a “fair” number of users incentivized across tasks, to further ensure fair training outcomes. To avoid excessive overhead during training, this client recruitment occurs before training begins [37]. In this section, we refer to potential clients as ‘users’, as they have not yet committed to the training. We model user $i$’s unwillingness towards training task $s$ by supposing $i$ will not train task $s$ unless they receive a payment $\geq c_{i,s}$. Following prior work [15], we suppose the $c_{i,s}$ are private and employ an auction that does not require users to explicitly report $c_{i,s}$; auctions require no prior information about users’ valuations, which may be difficult to obtain in practice. Users will submit bids $b_{i,s}$ indicating the payment they want for each task. Given the bids, the server will choose the auction winners (which users will participate in training) and each user’s payment. User $i$’s utility for task $s$ will be
@@
where $p_{i,s}$ is the payment received. We suppose that there is a shared budget $B$ across tasks. Since $B$ is unlikely to be sufficient to incentivize all users to perform all tasks, we seek to distribute the budget so as to ensure fair training across the tasks. We consider two types of “fairness:” budget-fairness, i.e., an equal budget for each task; and max-min fairness, distributing the budget so as to maximize the minimum take-up rate across tasks. Budget fairness seeks to equalize the resources (in terms of budget) across task, while max-min fairness seeks to make the training outcomes fair.
@@ V-A Fairness With Respect to the Same Budget across Tasks
$\max_{\textbf{x}_{s}\in\{0,1\},\textbf{p}}\ \sum_{s\in S}\left(\sum_{i\in I}x_{i,s}\right)$ (12)
@@
The decision variables for the server/ training coordinator are $\mathbf{x}_{s}$, a binary vector indicating whether each user is incentivized for task $s$, and $\mathbf{p}$, the set of payments to each user. The first constraint is the budget constraint for each task, and the second is that the user $i$ agrees to train task $s$ if $u_{i,s}\geq 0$.
@@
While this budget-fair auction is truthful, tasks disfavored by users (i.e., for which all users have high $c_{i,s}$) may still be able to incentivize few clients, leading to poor training performance. Thus, we next look at max-min fairness.
@@ V-B Fairness with Respect to Max-Min Task Take-Up Rates
$\max_{\textbf{x}_{s}\in\{0,1\},\textbf{p}}\min_{s\in S}\ (\sum_{i\in I}x_{i,s})\ \text{s.t.}\ \sum_{s\in S}\sum_{i\in I}p_{i,s}x_{i,s}\leq B,\,p_{i,s}\geq b_{i,s}$ (14)
As in (12)’s formulation, the decision variables are $\mathbf{x}_{s}$, and $\mathbf{p}$. The first constraint is the budget constraint, where the budget $B$ is coupled across the tasks $s\in S$. The second constraint is that users’ payments must exceed their bids. We first provide an algorithm to solve (14) in Algorithm 2, which greedily adds a user to each task until the budget $B$ is exhausted.
@@
5:$t=1$. Let $\underline{b}_{t,s}$ denote the $t$-th smallest bid.
@@
8:         $S_{s}\leftarrow S_{s}\cup t$; add user with minimum bid to $S_{s}$
9:         $B\leftarrow B-b_{t}$, update remaining budget
@@
Proof sketch. For an optimal max-min objective value $t$, we show by contradiction that all users with bids $\leq\underline{b}_{t}$ must be included in the optimal set. Algorithm 2 finds this solution. See the Technical Report [36] for the full proof. ∎.
While Algorithm 2 is optimal, it is not truthful: any user $i$ in the winning set for a task $s$ can easily increase its payment by increasing its bid $b_{i,s}>c_{i,s}$. Thus, we next design an auction method that, like Algorithm 2, proceeds in rounds to add an additional user to each task. We adapt the payment rule from the budget-fair auction in Section V-A to help ensure truthfulness. Our novel max-min fair auction is presented in Algorithm 3. The main idea, is to begin with a budget-fair auction for each task, and allocate users to tasks in rounds until a task runs out of budget (lines 4-9), as in Algorithm 2. We then re-allocate the budget across tasks so as to keep allocating users to all tasks, if possible (lines 11-16). If there is insufficient budget to do so, we can optionally perform a fractional allocation that uses up the remaining budget while remaining fair (lines 17-23). This ensures max-min fairness, and that the difference in number of users allocated across tasks is at most a fraction. In the learning process, this fraction can be modelled in terms of the time spent training tasks.
@@
1:Initialization: $B^{0}_{s}\leftarrow\frac{B}{s}$, equal initial budget $\forall$ tasks.
@@
5:         Let $\underline{b}_{i,s}$ denote the $i$-th biggest bid.
6:         if $\underline{b}_{i,s}\leq\frac{B}{i}$ then
@@
8:             and the Payment is updated $p_{i,s}\leftarrow\frac{B_{s}^{t}}{t}$.
@@
11:    if  $\exists\geq 1$ task s.t. the $i$-th user does not join then
12:         Let $S_{i-1}$ and $S_{i}$ denote the set of tasks with $i-1$ and $i$ users respectively.
13:         $A=\sum_{s\in S_{i-1}}[\underline{b}_{i,s}i-B_{s}^{i-1}]$
14:         $C=\sum_{s\in S_{i}}[B_{s}^{i-1}-\underline{b}_{i,s}i]$
@@
16:             Re-allocate amount $A$ from $S_{i}$ to $S_{i-1}$, obtaining $A$ from $s\in S_{i}$ via a waterfilling manner.
@@
18:             Fractionally allocate $C$ across $S_{i-1}$.
19:             $\underline{p}_{i,s}=\begin{cases}\frac{C}{|S_{i-1}|}\ &\text{if}\frac{C}{|S_{i-1}|}\leq\underline{b}_{i,s}\\ \underline{b}_{i,s}&\text{otherwise}\end{cases}$
20:             $\underline{x}_{i,s}=\begin{cases}\frac{C}{|S_{i-1}|\ \underline{b}_{i,s}}&\text{if}\frac{C}{|S_{i-1}|}\leq\underline{b}_{i,s}\\ 1&\text{otherwise}\end{cases}$
@@
Without the re-allocation across tasks, we would expect this auction to be truthful. Indeed, we show that such re-allocation rounds are the only instances that incentivize un-truthfulness:
@@
Proof Sketch: We characterize the untruthful cases (all during re-allocation rounds). The negation of the conditions a) and b) indicate that the auction does not end at other bids $\underline{b}_{j,s}$ smaller than the untruthful bid $\hat{b}_{i,s}$. We analyze all cases in detail in the Technical Report [36].
In practice, it would be difficult for users to exploit these conditions: all four conditions rely on the un-truthful bid comprising part of a re-allocation round. This heavily depends on other users’ bids, which would not be known. Moreover, since re-allocations are more likely to happen near the end of the auction, such un-truthful bids have a high risk of not being in the winning set at all. Next, we show that even if a user successfully submits such an un-truthful bid, it would have limited impact on the max-min fairness objective:
@@
Proof Sketch: The proof follows from analyzing each case in Theorem 8, as detailed in the Technical Report ( [36]). ∎
We now compare our max-min fair auction to the (truthful) budget-fair auction. Intuitively, we expect that it may lead to more fair outcomes. We can quantify this effect by studying the extreme case in which one task has no users, even if user bids from two tasks follow the same distribution:

### v2

@@ IV-B Convergence Guarantees
We firstly define $\overline{f}_{s}(\alpha)$ as:
$\overline{f}_{s}(\alpha)=\frac{f_{s}\left(\textbf{w}_{s}(t+1)\right)^{\alpha}}{\sum_{s^{\prime}}f_{s^{\prime}}\left(\textbf{w}_{s}(t+1)\right)^{\alpha}}.$ (28)
We rewrite
${\color[rgb]{0,0,0}\sum_{{\color[rgb]{0,0,0}Sel_{s}(t)}\in\mathbb{S}(t)}\sum_{k\in{\color[rgb]{0,0,0}Sel_{s}(t)}}B_{{\color[rgb]{0,0,0}Sel_{s}}}^{s}(\alpha)\left(\frac{p_{k,s}}{\sum_{k^{\prime}\in S(t)}p_{k’,s}}\right)^{2}}$ (29)
as $\sum_{j=1}^{K}\frac{1}{j}\left(\overline{f}_{s}(\alpha)\right)^{j}\left(1-\overline{f}_{s}(\alpha)\right)^{K-j}{K\choose j}.$ Since $\sum_{j=1}^{K}\left(\overline{f}_{s}(\alpha)\right)^{j}\left(1-\overline{f}_{s}(\alpha)\right)^{K-j}{K\choose j}=1$, we can view this as a weighted average of $1,\frac{1}{2},\ldots,\frac{1}{K}$. It then suffices to show that, as $\alpha$ increases, we place more weight on smaller terms. Note that
${\color[rgb]{0,0,0}\frac{d}{d\alpha}\left(\overline{f}_{s}(\alpha)^{j}\left(1-\overline{f}_{s}(\alpha)\right)^{K-j}\right)\propto\left(\frac{d\overline{f}_{s}(\alpha)}{\alpha}\right)\left(j-K\overline{f}_{s}(\alpha)\right)}$ (30)
It then suffices to show that $\frac{d\overline{f}_{s}(\alpha)}{\alpha}>0$; if so, the terms with larger $j$ will increase as $\alpha$ increases. Hence, $\frac{d\overline{f}_{s}(\alpha)}{\alpha}\propto\sum_{s^{\prime}=1}^{S}f_{s^{\prime}}^{\alpha}\left(\log f_{s}-\log f_{s^{\prime}}\right)>0$ as we have assumed task $s$ has the highest loss $f_{s}$. ∎
According to Corollary 5, as the fairness parameter $\alpha$ increases, our client-task allocation algorithm ensures that the worst-performing task makes more progress towards its training goal, thus helping to equalize the tasks’ performance and fulfilling the intuition of a higher $\alpha$ leading to more fair outcomes. The result further suggests that an alternative client-task allocation algorithm can explicitly enforce fairness across the upper-bounds in (4). However, this approach requires integer programming methods and estimating unknown constants like $\sigma^{2}$ and $L$ for each task, making it less practical than FedFairMMFL.
Finally, we provide the convergence bound for FedFairMMFL:
@@
With Assumptions 1-4, under a decaying learning rate $\eta_{t}=\frac{1}{\mu(t+\gamma)}$ and under our difficulty-aware client-task allocation algorithm FedFairMMFL, the error after $T$ iterations of federated learning for a specific task $s$ in multiple-model federated learning is
@@
Proof Sketch. We use the upper bound of $\mathbb{E}\left[\|\overline{\textbf{w}}^{(t+1)}_{s}-\textbf{w}_{s}^{*}\|^{2}\right]$ from Theorem 4, which incorporates our client-task allocation scheme. Next, we upper bound
${\color[rgb]{0,0,0}\Delta_{t+1}=\sum_{{\color[rgb]{0,0,0}Sel_{s}(t)}\in\mathbb{S}}B_{Sel_{s}}^{s}(\alpha)P^{k,2}_{Sel_{s}}}$ (31)
by $\frac{1}{2}$. Firstly, the inner summation $P^{k,2}_{Sel(t)}$ can be re-written as
$\sum_{k,k^{\prime}\in Sel_{s}(t),k\neq k^{\prime}}\frac{p_{k,s}}{(\sum_{k^{{}^{\prime\prime}}\in Sel_{s}(t)}p_{k^{\prime\prime},s})}\frac{p_{k^{\prime},s}}{(\sum_{k^{{}^{\prime\prime}}\in Sel_{s}(t)}p_{k^{\prime\prime},s})}$
$=\frac{(\sum_{k^{{}^{\prime\prime}}\in Sel_{s}(t)}p_{k^{\prime\prime},s})^{2}-\sum_{k\in Sel_{s}(t)}p_{k,s}^{2}}{2(\sum_{k^{{}^{\prime\prime}}\in Sel_{s}(t)}p_{k^{\prime\prime},s})^{2}}$
$\leq\frac{1}{2}.$
Secondly, the outer summation $\sum_{Sel_{s}(t)}B_{Sel_{s}}^{s}(\alpha(t))$ is a binomial distribution with binomial parameter $f_{s}^{\beta}(t)/{\sum_{s^{\prime}\in S}f_{s^{\prime}}^{\beta}(t)}$, and this sums to 1. Now we have a bound of $\Delta_{t+1}$ in terms of $\Delta_{t}$ and the learning rates $\eta_{t}$. Next, we use induction to bound $\Delta_{t+1}$ in terms of time independent variables. Finally, we use the L-smoothness of $F_{s}$ to obtain the result of this corollary. The full proof is available in our Technical Report [41].
As seen in the convergence bound, convergence to the optimum is impacted by the following parameters and effects:
First term: Our upper bound on $\mathbb{E}[f_{s}(\overline{\textbf{w}}^{(T)}_{s})]-f_{s}^{*}$ is quadratically proportional to the number of local epochs $\tau$ and the bound of the stochastic gradient $G$, and increases linearly with the bound of the variance of the stochastic gradient $\sigma^{2}$. Intuitively, as the number of local epochs or the stochastic gradient’s variance increase, convergence would be slower (i.e., our upper bound would be larger) due to greater client divergence and gradient noise, respectively. At the same time, our upper bound decreases in proportion to $\overline{\rho}^{\alpha}$, the lower bound on the selection skew. Intuitively, we can expect that as the selection skew increases, convergence will accelerate (i.e., our upper bound will be lower), as we have been selecting clients that were further from the global optimum with a higher probability. The convergence bound is also inversely proportional to $\mu^{2}$, which indicates that the higher the $\mu$-convexity, the faster the convergence.
Second term: Our upper bound is proportional to the square of $L$, the Lipschitz constant, as well as to $\Gamma_{s}$, the gap between the loss evaluated at the global optimum $\textbf{w}_{s}^{*}$ and that evaluated at client $k$’s optimal $w_{k,s}^{*}$ for task $s$, summed over all clients. Intuitively, if the data distribution is more non-iid (meaning $\Gamma_{s}$ is larger, convergence will be slower, as shown by an increase in our upper bound. Convergence is inversely proportional to the square of $\mu$, which as for the first term indicates that the higher the $\mu$-convexity, the faster the convergence.
Third term: As the discrepancy between the initial point of training and the global optimum for task $s$ increases, the model requires more iterations to approach the optimum.
Fourth term: Similar to the second term, convergence is proportional to $L$, the Lipschitz constant, as well as to $\Gamma_{s}$, the gap between the loss evaluated at the global optimal $\textbf{w}_{s}^{*}$ and the loss evaluated at client $k$’s optimal $w_{k,s}^{*}$ for task $s$, summed over all clients. Convergence is inversely proportional to $\mu$, which indicates that the higher the $\mu$-convexity, the faster the convergence. The term $(\frac{\tilde{\rho}^{\alpha}}{\overline{\rho}^{\alpha}}-1)$ indicates the gap between the upper bound and lower bound of the selection skew, and hence the bias of the selection skew. As the bias increases, convergence slows down, i.e., our upper bound is larger, since the selection is biased to certain clients.
@@ V Client Incentive-Aware Fairness across Tasks
In this section, we account for the fact that clients may have different willingness to train each model. Prior incentive-aware FL work [36, 40, 21] proposed auctions to incentivize clients to join FL training. These works aimed to maximize the total number of users performing either one (or many) tasks. However, in the multiple model case, when many users favor a particular task over others, the performance of the disfavored task can suffer, as fewer users may be incentivized to perform this task. Here, we would like to ensure a “fair” number of users incentivized across tasks, to further ensure fair training outcomes. To avoid excessive overhead during training, the client recruitment auctions in this section occur before training begins [42]. In this section, we refer to potential clients as ‘users’, as they have not yet committed to the training.
We model user $i$’s unwillingness towards training task $s$ by supposing $i$ will not train task $s$ unless they receive a payment $\geq c_{i,s}$. Following prior work [21], we suppose the $c_{i,s}$ are private and employ an auction that does not require users to explicitly report $c_{i,s}$; auctions require no prior information about users’ valuations, which may be difficult to obtain in practice. Users will submit bids $b_{i,s}$ indicating the payment they want for each task. Given the bids, the server will choose the auction winners who will participate in MMFL training (second part of pipeline in Fig. 1), and each user’s payment. User $i$’s utility for task $s$ will be
@@
where $p_{i,s}$ is the payment received.
We suppose that there is a shared budget $B$ across tasks. Since $B$ is unlikely to be sufficient to incentivize all users to perform all tasks, we seek to distribute the budget so as to ensure fair training across the tasks. We consider two types of “fairness:” budget-fairness, i.e., an equal budget for each task; and max-min fairness, distributing the budget so as to maximize the minimum take-up rate across tasks. Budget fairness seeks to equalize the resources (in terms of budget) across task, while max-min fairness seeks to make the training outcomes fair.
@@ V-A Fairness With Respect to the Same Budget across Tasks
$\max_{\textbf{x}_{s}\in\{0,1\},{\color[rgb]{0,0,0}\textbf{P}}}\ \sum_{s\in S}\left(\sum_{i\in I}x_{i,s}\right)$ (33)
@@
$\,p_{i,s}\geq b_{i,s}\mathbbm{1}_{\{x_{i,s}\geq 0\}},\ \forall i.$ (35)
The decision variables for the server/ training coordinator are $\mathbf{x}_{s}$, a binary vector indicating whether each user is incentivized for task $s$, and $\mathbf{P}=\{...,p_{i,s},...\}$, the set of payments to each user for each task. The first constraint is the budget constraint for each task, and the second is that the user $i$ agrees to train task $s$ if $u_{i,s}\geq 0$.
@@
While this budget-fair auction is truthful, tasks disfavored by users (i.e., for which all users have high $c_{i,s}$) may only be able to incentivize few clients, leading to poor training performance. Thus, we next look at max-min fairness.
@@ V-B Fairness with Respect to Max-Min Task Take-Up Rates
$\max_{\textbf{x}_{s}\in\{0,1\},{\color[rgb]{0,0,0}\textbf{P}}}\min_{s\in S}\ (\sum_{i\in I}x_{i,s})$ (36)
[…]
$p_{i,s}\geq b_{i,s}$ (38)
As in (33)’s formulation, the decision variables are $\mathbf{x}_{s}$, and ${\color[rgb]{0,0,0}\mathbf{P}}$. The first constraint is the budget constraint, where the budget $B$ is coupled across the tasks $s\in S$. The second constraint is that users’ payments must exceed their bids. We first provide an algorithm to solve (36) in Algorithm 2, which loops across tasks and greedily adds a user to each task until the budget $B$ is exhausted.
@@
5:$t=1$. Let $\underline{b}_{t,s}$ denote the $t$-th smallest bid for task $s$, and $\underline{u}_{t,s}$ denote the corresponding user.
@@
8:         $S_{s}\leftarrow S_{s}\cup{\color[rgb]{0,0,0}\underline{u}_{t,s}}$; add user with minimum bid to $S_{s}$
9:         $B\leftarrow B-{\color[rgb]{0,0,0}\underline{b}_{t,s}}$ update remaining budget
@@
Proof sketch. Let $\underline{b}_{t,s}$ denote the $t$-th smallest bid for task $s$, and $\underline{u}_{t,s}$ denote the corresponding user. For an optimal max-min objective value $t$, we show by contradiction that all users with bids $\leq{\color[rgb]{0,0,0}\underline{b}_{t,s}}$ must be included in the optimal set. Algorithm 2 finds this solution. See the Technical Report [41] for the full proof. ∎.
While Algorithm 2 is optimal, it is not truthful: any user $i$ in the winning set for a task $s$ can easily increase its payment by increasing its bid $b_{i,s}>c_{i,s}$. Thus, we next design an auction method that, like Algorithm 2, proceeds in rounds to add an additional user to each task. We adapt the payment rule from the budget-fair auction in Section V-A to help ensure truthfulness. Our novel max-min fair auction is presented in Algorithm 3. The main idea, is to begin with a budget-fair auction for each task, and allocate users to tasks in rounds until a task runs out of budget (lines 4-9), as in Algorithm 2. Once a task runs out of budget, we then re-allocate the budget across tasks so as to keep allocating users to all tasks, if possible (lines 11-16). If there is insufficient budget to do so, we can optionally perform a fractional allocation that uses up the remaining budget while remaining fair (lines 17-23). This ensures max-min fairness, and that the difference in number of users allocated across tasks is at most a fraction. In the learning process, this fraction can be modelled in terms of the time spent training tasks. As mentioned earlier, the auction takes place before training, to avoid additional overhead.
@@
1:Initialization: $B^{0}_{s}\leftarrow\frac{B}{s}$, equal initial budget $\forall$ tasks, where $B^{t}_{s}$ is the budget allocated to task $s$ at time $t$.
@@
5:          Let $\underline{b}_{t,s}$ denote the $t$-th biggest bid, for task $s$, and $\underline{u}_{t,s}$ denote the corresponding user.
6:         if $\underline{b}_{t,s}\leq\frac{B}{t}$ then
@@
8:             and the Payment is updated $\underline{p}_{t,s}\leftarrow\frac{B_{s}^{t}}{t}$.
@@
11:    if  $\exists\geq 1$ task s.t. the $t$-th user does not join then
12:         Let $S_{t-1}$ and $S_{t}$ denote the set of tasks with $t-1$ and $t$ users respectively.
13:         $A=\sum_{s\in S_{t-1}}[\underline{b}_{t,s}t-B_{s}^{t-1}]$
14:         $C=\sum_{s\in S_{t}}[B_{s}^{t-1}-\underline{b}_{t,s}t]$
@@
16:             Re-allocate amount $A$ from $S_{t}$ to $S_{t-1}$, obtaining $A$ from $s\in S_{t}$ via a waterfilling manner.
@@
18:             Fractionally allocate $C$ across $S_{t-1}$.
19:             $\underline{p}_{t,s}=\begin{cases}\frac{C}{|S_{t-1}|}\ &\text{if}\frac{C}{|S_{t-1}|}\leq\underline{b}_{t,s}\\ \underline{b}_{t,s}&\text{otherwise}\end{cases}$
20:             $\underline{x}_{t,s}=\begin{cases}\frac{C}{|S_{t-1}|\ \underline{b}_{t,s}}&\text{if}\frac{C}{|S_{t-1}|}\leq\underline{b}_{t,s}\\ 1&\text{otherwise}\end{cases}$
@@
Without the re-allocation across tasks, we would expect this auction to be truthful. Indeed, in Theorem 8, we show that such re-allocation rounds are the only instances that incentivize un-truthfulness. In practice, it would be difficult for users to exploit these conditions: all four conditions rely on the un-truthful bid comprising part of a re-allocation round. This heavily depends on other users’ bids, which would not be known. Moreover, since re-allocations are more likely to happen near the end of the auction, such un-truthful bids have a high possibility of not being in the winning set at all.
@@
Proof.
We characterize the untruthful cases (all during re-allocation rounds). The negation of the conditions a) and b) indicate that the auction does not end at other bids $\underline{b}_{j,s}$ smaller than the untruthful bid $\hat{b}_{i,s}$.
Case 1: c) re-allocation required for the un-truthful bid and d) there is sufficient budget in $S_{t}$ for re-allocation. Note that the truthful bid either entered in an earlier round, or also received re-allocation.
Case 2: For both the truthful and un-truthful bid, budget re-allocation across task needs to happen (e and g), but there are insufficient funds (f and h). Fractional allocation happens, and this un-truthful strategy dominates when $\min(b_{i,s},\frac{C}{|S_{t-1}|})<\min(\hat{b}_{i,s},\frac{C}{|S_{t-1}|})$ (i).
Case 3: The truthful bid enters the auction at an earlier round $l$ (k), there is insufficient budget for the untruthful bid (l), and insufficient budget for re-allocation (j). Fractional allocation happens, and this untruthful strategy dominates when either $\hat{b}_{i,s}<\frac{C}{|S_{t-1}|}$ or $\frac{C}{|S_{t-1}|}>\frac{B^{l}_{s}}{l}$ (m).
Case 4: Re-allocation is required for both the truthful and untruthful bid (n and p). There is sufficient budget for re-allocation for the truthful bid (o), unlike the untruthful bid (q). This untruthful strategy dominates if $b_{i,s}<\min(\frac{C}{|S_{t-1}|},\hat{b}_{i,s})$. ∎
Next, we show that even if a user successfully submits an un-truthful bid, it would have limited impact on the max-min fairness objective:
@@
Proof.
The proof follows from analyzing each case in Theorem 8.
Case 1b: Case 1 can be divided into two cases: a) the truthful bid ($T$) entering in an earlier round, and b) the truthful bid receiving re-allocation. For both cases what is common is that the untruthful bid ($U$) receives re-allocation. For case b) let $a_{1}$ and $a_{2}$ be bids greater than $U$ for task $s$ ($a_{1}$ need not be the next greatest bid after $U$, but $a_{2}$ is the next greatest bid after $a_{1}$). Proving by contradiction we assume that both $a_{1},a_{2}$ are winners under the truthful setting and non-winners in the non-truthful setting. The situation $T+a_{1}<b_{r},U+a_{1}>b_{r}$ is possible, causing the fractional round to come one round earlier when the un-truthful bid is made. (Where $b_{r}$ is the remaining budget after taking out the budget re-allocated to other bids in other tasks, and re-allocated to other bids in task $s$ between $U$ and $a_{1}$.) The situation $T+a_{1}+a_{2}<b_{r},U+a_{1}>b_{r}$ is not possible, as $a_{2}$ is greater than $U$. Hence the fractional round comes at most one round earlier when the un-truthful bid is made.
Case 4: The same reasoning as that above works for Case 4, where $a_{1}$ in the untruthful case is replaced by $U$, as $U$ is the bid in the fractional allocation round. WLOG $U$ comes directly after $o_{3}$. We have [$T+o_{1}+o_{2}+o_{3}<b_{r},o_{1}+o_{2}+o_{3}+U>b_{r}$] possible but [$T+o_{1}+o_{2}+o_{3}+a_{2}<b_{r},o_{1}+o_{2}+o_{3}+U>b_{r}$] impossible as $U<a_{2}$. Thus the gap in max-min objective in the truthful vs untruthful setting is at most one.
Case 1a: We let $o_{1},o_{2}$, etc.. be the bids next in order after $T$. Supposing for the truthful setting we have $[T<\frac{B}{t},o_{1}<\frac{B}{t+1},o_{2}<\frac{B}{t+2}]$ and for the untruthful setting we have $[o_{1}<\frac{B}{t},o_{2}<\frac{B}{t+1},o_{3}\nless\frac{B}{t+2}]$ (where $U$ is larger than $o_{3}$) - this means that in the untruthful setting $o_{3}$ is the first round where re-allocation occurs. Going back to the truthful setting, we will be unable to have "$o_{3}<\frac{B}{t+3}"$, because $\frac{B}{t+3}<\frac{B}{t+2}$ and we already have $o_{3}\nless\frac{B}{t+2}$ from the untruthful setting. Hence the re-allocation round occurs at most one round earlier under the un-truthful setting, given that all other bids are fixed. Next, to show that the fractional round comes at most one round earlier, we use a similar reasoning as the above cases: $[o_{3}+o_{4}+a_{1}<b_{r},o_{3}+o_{4}+U+a_{1}>b_{r}]$ is possible, while $[o_{3}+o_{4}+a_{1}+a_{2}<b_{r},o_{3}+o_{4}+U+a_{1}>b_{r}]$ is not possible since $a_{2}>U$.
Case 3: The link between Case 3 and Case 1a, is analogous to the link between Case 4 and Case 1b. Hence, a similar reasoning to that of 1a works for Case 3, where $a_{1}$ in the untruthful case will be replaced by $U$, as $U$ is the bid in the fractional allocation round. WLOG $U$ comes directly after $o_{5}$. We thus have $[o_{3}+o_{4}+o_{5}<b_{r},o_{3}+o_{4}+o_{5}+U>b_{r}]$ being possible while $[o_{3}+o_{4}+o_{5}+a_{2}<b_{r},o_{3}+o_{4}+o_{5}+U>b_{r}]$ being impossible as $U<a_{2}$. This means that the fractional round comes at most one round earlier, in the untruthful setting.
Case 2: For case 2, both the truthful and un-truthful bids lead to a fractional re-allocation at the $t$th round, after which the auction ends. Fractional allocation happens at the same round for both, because there is only one round of fractional allocation under our auction, and Otherwise, a bid between $T$ and $U$ would have entered, and $U$ will not be in the winning set. Thus the gap in max-min objective in the truthful vs untruthful setting is at most one.
When the fractional round comes at most one round earlier (as in some of the above cases), difference in max-min objective can be bounded by 2.
We now compare our max-min fair auction to the (truthful) budget-fair auction in Section V-A. Intuitively, we expect that it may lead to more fair outcomes. We can quantify this effect by comparing the outcomes of the two auctions, for the extreme case in which one task has no users, even if user bids from two tasks follow the same distribution: